Please help me answer this equation +-20=-4x-6x+
I tried it but I couldnt get it
This is what I did
-20=-4x-6x
-20=-10x divided both by 10 and got
-10=-x
10=x But when i tried to check it it didnt work
-20=-4*10-6*-10
-20=-40-60
-20=100

Found 2 solutions by richwmiller, algebrapro18:
Answer by richwmiller(17219)

You added -20+10=-10
You didn't divide -20/10
-20=-4x-6x
-20=-10x
2=x
-20=-4x-6x
-20=-4*2-6*2
-20=-8-12
-20=-20

Answer by algebrapro18(249)

you did the right steps but -20/-10 is 2 not 10.